# Simple Fermionic backflow states via a systematically improvable tensor decomposition

### v1

@@ III.2 Water molecule
Figure 3: Relative ground state correlation energy error (compared to exact diagonalization) for CPD backflow ansätze ($\Psi^{CPD}$) on the water molecule (6-31G basis, equilibrium geometry as specified in Ref. [4]) as the number of configurational samples in each Markov chain is increased. Two support dimensions corresponding to $M=1$ and $M=4$ are shown. Gaussian process state (GPS) and restricted Boltzmann machine (RBM) comparison energies are taken from Refs. [15] and [4], respectively, while the CCSD energy is calculated with PySCF [61, 62].
@@
The results of the CPD backflow state for different number of configurational samples in this system are shown in Fig. 3. Once again, we find that increasing the number of samples can affect the accuracy more significantly than increasing the complexity of the model, showing results going from $M=1$ ($6.7k$ parameters) to $M=4$ ($27k$ parameters). Nevertheless, we outperform the accuracy of the native RBM state ($728$ parameters) with up to $10^{6}$ samples [4], as well as the GPS multiplied by a symmetry-broken Pfaffian state ($585$ parameters) with about $10^{4}$ samples [15], approaching the accuracy of coupled-cluster with singles and doubles level of theory which excels for weakly correlated equilibrium molecular systems such as this. While we still don’t quite reach this level (and the significant ‘chemical accuracy’ hurdle – albeit defined with respect to the finite basis set energy), to the best of our knowledge, these results represent the state-of-the-art for an NQS-like variational ansatz for this system, reaching a relative correlation energy error of $\approx 2\%$ with support dimension $M=4$ and using a comparatively small number of $\mathcal{O}[10^{4}]$ samples.
@@ III.3 Towards hydrogen materials
Figure 4: Ground state energy per atom of a $6\times 6$ square hydrogen lattice in a STO-6G basis for increasing lattice constants. Shown (top) are RHF (orange) and UCCSD energies (violet), while the horizontal grey line indicates the exact energy of the fully dissociated limit in this basis. The CPD backflow ansatz ($\Psi^{CPD}$) with support dimension $M=1$ (solid black line) is compared to a GPS ansatz augmented with a Slater determinant ($\Psi^{GPS}\times|\Phi\rangle$) with support dimension $M=72$ [15] (dotted black line), as well as energies obtained with DMRG (red dots and line) using a bond dimension of $M=1024$. A further CPD backflow ansatz with support dimension $M=1$ and a truncation in the backflow subspace to $K=5$ orbitals is shown as a dashed black line. The backflow CPD (GPS augmented by Slater determinant) VMC ansätze were optimized with $4096$ ($10,000$) samples, while DMRG results were obtained using the block2 package [75, 76]. The bottom panel shows energy difference relative to DMRG of the CPD backflow ansätze and the GPS times Slater determinant ansatz.
In Figure 4, we report the ground state energy per atom obtained from the CPD backflow ansätze (with and without the rank and range truncation of Sec. II.2 for the backflow transform to $K=5$ sites) for a $6\times 6$ hydrogen lattice in a minimal basis (STO-6G). We consider both compact (lower effective $U/t$) and extended (higher effective $U/t$) lattice structures by varying the interatomic distances all the way to essentially dissociated non-interacting hydrogen atoms. We compare our results to energies obtained with restricted Hartree-Fock (RHF) and unrestricted coupled-cluster with single and double excitations (UCCSD), as well as an efficient ab initio implementation of density matrix renormalization group (DMRG) going up to bond dimension of $1024$ in a fully spin-adapted basis implemented in the block2 package [75, 76]. For an additional comparison between contrasting approaches to Fermionic variational wave functions, we also include the results obtained from a GPS ansatz with support dimension $M=72$ acting as a Jastrow in front of a co-optimized Slater determinant [15], to compare the CPD backflow to this approach. We optimize the CPD backflow and GPS multiplied by Slater determinant ansätze in a Boys localized basis for the orbitals, whereas for the DMRG results we rely on a split-localized basis, in which occupied and virtual orbitals are localized separately. We found this choice to give the most consistent results for DMRG across the range of geometries studied.
@@
Table 1: Dissociation energy ($D_{e}$), equilibrium bond length ($r_{e}$), harmonic ($\omega_{e}$) and anharmonic ($\omega_{e}\chi_{e}$) frequencies in wavenumbers for the $6\times 6$ hydrogen lattice obtained by fitting a Morse potential to data from different quantum chemistry and variational methods described in this work.
@@ III.3.1 Spin-spin correlations
Figure 5: Two-point instantaneous spin-spin correlation function for the ground state of the $6\times 6$ hydrogen lattice between atoms at two different interatomic distances $d=1.2$ and $d=3.0$. Correlation function is given as a function of normalized radial interatomic distance between atoms in the correlator, showing nearest, next-nearest and next-next-nearest magnetic correlations. The correlators were computed from the optimized CPD backflow ansatz with $M=1$.
@@ IV Scaling
As illustrated in the results above, the CPD backflow ansatz performs well on small Fermionic systems, but further developments for scaling to significantly larger systems are still required for this to become a clearly competitive method for the wider electronic structure community. Simplifying the scaling to assume a general growth of both the basis and electron number such that $N\sim L$ and assuming $M$ and $K$ are independent of system size, the parameters grow with system size as $N_{P}\sim\mathcal{O}[N^{3}]$ for the full ansatz, with the subspace truncation of the backflow reducing this asymptotically to $N_{P}\sim\mathcal{O}[N^{2}]$ (see Sec. II.2). The fast updating of backflow orbitals also enables the evaluation of the wave function log-amplitudes to be performed in $\mathcal{O}[MN^{2}+N^{3}]$ (regardless of whether a backflow truncation is applied). However, we find in practice the determinant evaluation has a significantly smaller prefactor than the construction of the orbitals, so that the dominant scaling is rather $\mathcal{O}[N^{2}]$ for small to medium-sized systems. Given that the number of terms in general second quantized ab initio hamiltonians scales as $\mathcal{O}[N^{4}]$, the resulting scaling of the local energy evaluation is then $\mathcal{O}[N^{7}]$ for the CPD backflow state in the asymptotic limit, or $\mathcal{O}[N^{6}]$ for small to medium-sized systems. While this should be competitive with accurate quantum chemical methods such as coupled-cluster, it is clear that the prefactor is significantly larger.
Rather than just the local energy evaluation, we should also consider the computational scaling for the update of the parameters. For larger numbers of parameters (such as the $\sim 187,000$ of the hydrogen lattice above), their update used to be the main bottleneck for VMC large-scale ansätze when using the original SR algorithm [56]. For a model with $N_{P}$ parameters, SR would scale as $\mathcal{O}(N_{P}^{3})$, since it involves inverting the $N_{P}\times N_{P}$ quantum geometric tensor matrix. This is not the case with the kernel formulation of SR recently introduced in Ref. [57], which takes advantage of a simple linear algebra identity to reduce the dimension of the matrix that is inverted in the SR algorithm from $N_{P}$ to $N_{S}$, i.e. the number of samples used during the optimization. When $N_{P}\gg N_{S}$, as in large-scale models, the scaling of the parameter update becomes $\mathcal{O}(N_{S}^{2}N_{P}+N_{S}^{3})$, i.e. linear in the number of parameters. Thus, the evaluation of the local energy remains the computational bottleneck of the algorithm in the case of ab initio systems.
@@
Figure 6: Mean VMC step runtime for the CPD backflow ansatz as a function of the number of atoms in a hydrogen chain with fixed inter-atomic distances ($1.68$) in a STO-6G basis. The support dimension is $M=1$ and the sample size is $N_{S}=128$. The blue points are obtained with the full Hamiltonian, while the green and orange points are obtained with the local energy evaluated with a Hamiltonian pruned with a threshold of $10^{-9}$ and $10^{-5}E_{h}$. The dashed lines show the observed scaling for the largest system sizes.
@@ V Conclusions and outlook
In this work, we introduce a general and simple ansatz suitable for ab initio fermions, based on a systematically improvable tensor rank decomposition of a general backflow form. This systematically builds configuration-dependent orbitals of a single antisymmetric Slater determinant in second quantization, directly encoding non-trivial $N$-body electron-electron correlations with a parameter scaling of $\mathcal{O}[N^{2-3}]$. We have shown that the ansatz can achieve competitive accuracy on small Fermionic systems, such as the Fermi-Hubbard model and the water molecule, and that it can be used to model larger strongly correlated lattices of ab initio hydrogen atoms with an accuracy comparable to state-of-the-art DMRG techniques. Finally, we have discussed the scalability of the ansatz and shown that we can affect various reductions in a practical fashion to demonstrate $\mathcal{O}(N^{4})$ scaling on medium to large ab initio systems. We are working on further improvements in the accuracy and efficiency of the ansatz, as well as taking advantage of the benefits of working in a second quantized formalism to integrate with multiscale methods and quantum embedding methodologies to provide a practical route in the modelling of truly extended systems within this CPD backflow framework [80, 80].
@@
The code for this project is implemented as part of the freely available GPSKet plugin for NetKet [59, 60]. It is made available, together with configurations files to reproduce the figures in the paper, at https://github.com/BoothGroup/GPSKet/tree/master/scripts/cpd-backflow.

### v2

@@ III.2 Water molecule
Figure 3: Systematic improvability of the CPD backflow ansatz on the water molecule with respect to both model complexity ($M$) and stochastic samples ($N_{S}$). Relative ground state correlation energy error (compared to exact diagonalization) for the CPD backflow ansatz ($\Psi^{CPD}$) on the water molecule (6-31G basis, equilibrium geometry as specified in Ref. [4]) as a function of (a) support dimension $M$ and (b) number of configurational samples $N_{S}$. We also extrapolate these results to infinite $M$ or $N_{S}$ to provide the values in their infinite limit as shown by diamonds.
The results in Fig. 3(a) show that the CPD backflow ansatz formulated in this local basis exhibits a strikingly systematic improvability, with the error decreasing inversely with the support dimension of the model, $M$. We can use this empirical scaling to extrapolate the results to the infinite support dimension limit, which results in a relative correlation energy error of below $2\%$, for the ansatz optimized with $\mathcal{O}[10^{4}]$ configurational samples. At infinite $M$ the model is complete, and the error therefore must arise from the incomplete optimization of the finite-$M$ models. We therefore also consider the improvability in $M$ for two different numbers of configurational samples, $N_{S}$ in the Markov chains used for each optimization step. The $1/M$ decay of the error is clearly seen in both of these sample sizes, with the extrapolated model result decreasing towards exactness for increasing $N_{S}$.
We analyse this trend more systematically in Fig. 3(b), where we show the convergence in $N_{S}$ for two different model complexity parameters $M$, showing a relatively robust $N_{S}^{-\frac{1}{2}}$ scaling in the error. This indicates that doubling the support dimension has a similar effect on reducing the error as quadrupling $N_{S}$. This robustness and reliability in the error reduction is to be expected with increasing $M$, but is more surprising with increasing $N_{S}$. It indicates that the noise introduced into the sampling at finite $N_{S}$ values is not simply changing the variance in the resulting energy, or indeed resulting in different optimized states due to convergence to different local minima in the landscape (where we would expect a wider scatter of optimized energies). Instead, the robustness and systematic trend in the results indicates that $N_{S}$ is controlling the intrinsic error of the optimization of the state in a more systematic fashion. This could potentially arise from non-linear steps in the optimization protocol, and is something which requires further scrutiny going forwards.
@@
Comparing the accuracies obtained to previous state-of-the-art results in Figure 4, we find that the CPD backflow state with $M=1$ ($6.7k$ parameters) already outperforms both the Gaussian process state augmented by a symmetry-broken Pfaffian ($585$ parameters) [15] and the restricted Boltzmann machine NQS state ($728$ parameters) [4] when optimized with $\mathcal{O}[10^{3}]$ configurational samples. The accuracy is further improved when larger support dimensions and sample sizes are considered, with the CPD model at $M=4$ ($27k$ parameters) and $N_{S}\sim\mathcal{O}[10^{4}]$ outperforming the best NQS by $2\%$ in the relative correlation energy error. While we still don’t quite reach the level of accuracy of coupled cluster methods with singles and doubles (and the significant ‘chemical accuracy’ hurdle – albeit defined with respect to the finite basis set energy), to the best of our knowledge, these results represent the state-of-the-art for an NQS-like variational ansatz for this system.
Figure 4: Performance of the CPD backflow and other models on the water molecule. Relative ground state correlation energy error (compared to exact diagonalization) for CPD backflow ansätze ($\Psi^{CPD}$) on the water molecule (6-31G basis, equilibrium geometry as specified in Ref. [4]) as the number of configurational samples $N_{S}$ in each Markov chain is increased. Two support dimensions corresponding to $M=1$ (blue) and $M=4$ (green) are shown. Comparison energies for the Gaussian process state augmented by a symmetry-broken Pfaffian ($\Psi^{GPS}\times\left|\text{Pf}\right\rangle$, dash-dotted line) and restricted Boltzmann machine NQS ($\Psi^{RBM}$, dotted and dashed lines) are taken from Refs. [15] and [4], respectively, while the CCSD energy (solid line) is calculated with PySCF [65, 66].
@@ III.3 Towards hydrogen materials
Figure 5: Performance of the CPD backflow and other methods on the hydrogen lattice. a Ground state energy per atom of a $6\times 6$ square hydrogen lattice in a STO-6G basis for increasing lattice constants. Shown are RHF (orange) and UCCSD energies (violet), while the horizontal grey line indicates the exact energy of the fully dissociated limit in this basis. The CPD backflow ansatz ($\Psi^{CPD}$) with support dimension $M=1$ (solid black line) is compared to a GPS ansatz augmented with a Slater determinant ($\Psi^{GPS}\times|\Phi\rangle$) with support dimension $M=72$ [15] (dotted black line), as well as energies obtained with DMRG (red dots and line) using a bond dimension of $M=1024$. A further CPD backflow ansatz with support dimension $M=1$ and a truncation in the backflow subspace to $K=5$ orbitals is shown as a dashed black line. The backflow CPD (GPS augmented by Slater determinant) VMC ansätze were optimized with $4096$ ($10,000$) samples, while DMRG results were obtained using the block2 package [80, 81]. b Energy difference per atom relative to DMRG of the CPD backflow ansätze and the GPS times Slater determinant ansatz. To aid the comparison, we have removed the outlier DMRG energy at $2.5$.
In Figure 5, we report the ground state energy per atom obtained from the CPD backflow ansätze (with and without the rank and range truncation of Sec. II.3 for the backflow transform to $K=5$ sites) for a $6\times 6$ hydrogen lattice in a minimal basis (STO-6G) with open boundary conditions. The choice of $K=5$ for the range cutoff of the backflow ensures that the local symmetries of quantum fluctuations about each atomic site are preserved, as discussed in Sec. II.3. We consider both compact (lower effective $U/t$) and extended (higher effective $U/t$) lattice structures by varying the interatomic distances all the way to essentially dissociated non-interacting hydrogen atoms. We compare our results to energies obtained with restricted Hartree-Fock (RHF) and unrestricted coupled-cluster with single and double excitations (UCCSD), as well as an efficient ab initio implementation of density matrix renormalization group (DMRG) going up to bond dimension of $1024$ in a fully spin-adapted basis implemented in the block2 package [80, 81]. For an additional comparison between contrasting approaches to Fermionic variational wave functions, we also include the results obtained from a GPS ansatz with support dimension $M=72$ acting as a Jastrow in front of a co-optimized Slater determinant [15], to compare the CPD backflow to this approach. We optimize the CPD backflow and GPS multiplied by Slater determinant ansätze in a Boys localized basis for the orbitals, whereas for the DMRG results we rely on a split-localized basis, in which occupied and virtual orbitals are localized separately. We found this choice to give the most consistent results for DMRG across the range of geometries studied. Each optimization of the CPD backflow wave functions took $\approx 250$ GPU hours across 4 Nvidia A100 devices, whereas the DMRG runs took a total of $\approx 500$ CPU hours on an Intel(R) Core(TM) i9 device.
@@
Table 1: Physical properties of the hydrogen lattice as obtained by different methods. Dissociation energy ($D_{e}$), equilibrium bond length ($r_{e}$), harmonic ($\omega_{e}$) and anharmonic ($\omega_{e}\chi_{e}$) frequencies in wavenumbers for the $6\times 6$ hydrogen lattice obtained by fitting a Morse potential to data from different quantum chemistry and variational methods described in this work.
@@ III.3.1 Spin-spin correlations
Figure 6: Spin-spin correlations in the hydrogen lattice. Two-point instantaneous spin-spin correlation function for the ground state of the $6\times 6$ hydrogen lattice between atoms at two different interatomic distances $d=1.2$ (blue) and $d=3.0$ (green). Correlation function is given as a function of normalized radial interatomic distance between atoms in the correlator, showing nearest, next-nearest and next-next-nearest magnetic correlations. The correlators were computed from the optimized CPD backflow ansatz with $M=1$.
@@ IV Scaling
As illustrated in the results above, the CPD backflow ansatz performs well on small Fermionic systems, but further developments for scaling to significantly larger systems are still required for this to become a clearly competitive method for the wider electronic structure community. Simplifying the scaling to assume a general growth of both the basis and electron number such that $N\sim L$ and assuming $M$ and $K$ are independent of system size, the parameters grow with system size as $N_{P}\sim\mathcal{O}[N^{3}]$ for the full ansatz, with the subspace truncation of the backflow reducing this asymptotically to $N_{P}\sim\mathcal{O}[N^{2}]$ (see Sec. II.3). The fast updating of backflow orbitals also enables the evaluation of the wave function log-amplitudes to be performed in $\mathcal{O}[MN^{2}+N^{3}]$ (regardless of whether a backflow truncation is applied). However, we find in practice the determinant evaluation has a significantly smaller prefactor than the construction of the orbitals, so that the dominant scaling is rather $\mathcal{O}[N^{2}]$ for small to medium-sized systems. Given that the number of terms in general second quantized ab initio hamiltonians scales as $\mathcal{O}[N^{4}]$, the resulting scaling of the local energy evaluation is then $\mathcal{O}[N^{7}]$ for the CPD backflow state in the asymptotic limit, or $\mathcal{O}[N^{6}]$ for small to medium-sized systems. While this should be competitive with accurate quantum chemical methods such as coupled-cluster, it is clear that the prefactor is significantly larger.
Rather than just the local energy evaluation, we should also consider the computational scaling for the update of the parameters. For larger numbers of parameters (such as the $\sim 187,000$ of the hydrogen lattice above), their update used to be the main bottleneck for VMC large-scale ansätze when using the original SR algorithm [60]. For a model with $N_{P}$ parameters, SR would scale as $\mathcal{O}(N_{P}^{3})$, since it involves inverting the $N_{P}\times N_{P}$ quantum geometric tensor matrix. This is not the case for recently introduced alternative formulations of SR, such as minimum-step SR [82], the kernel formulation of SR [61] or SPRING [71]. In particular for minimum-step SR and the kernel formulation of SR simple linear algebra identities were used to reduce the dimension of the matrix that is inverted in the SR algorithm from $N_{P}$ tos $N_{S}$, i.e. the number of samples used during the optimization. When $N_{P}\gg N_{S}$, as in large-scale models, the scaling of the parameter update becomes $\mathcal{O}(N_{S}^{2}N_{P}+N_{S}^{3})$, i.e. linear in the number of parameters. Thus, the evaluation of the local energy remains the computational bottleneck of the algorithm in the case of ab initio systems.
@@
Figure 7: Scalability of the CPD backflow ansatz. Mean VMC step runtime for the CPD backflow ansatz as a function of the number of atoms in a hydrogen chain with fixed inter-atomic distances ($1.68$) in a STO-6G basis. The support dimension is $M=1$ and the sample size is $N_{S}=128$. The blue points are obtained with the full Hamiltonian, while the green and orange points are obtained with the local energy evaluated with a Hamiltonian pruned with a threshold of $10^{-9}$ and $10^{-5}E_{h}$. The dashed lines show the observed scaling for the largest system sizes.
@@ V Conclusions and outlook
In this work, we introduce a general and simple ansatz suitable for ab initio fermions, based on a systematically improvable tensor rank decomposition of a general backflow form. This systematically builds configuration-dependent orbitals of a single antisymmetric Slater determinant in second quantization, directly encoding non-trivial $N$-body electron-electron correlations with a parameter scaling of $\mathcal{O}[N^{2-3}]$. We have shown that the ansatz can achieve competitive accuracy on small Fermionic systems, such as the Fermi-Hubbard model and the water molecule, and that it can be used to model larger strongly correlated lattices of ab initio hydrogen atoms with an accuracy comparable to state-of-the-art DMRG techniques. Finally, we have discussed the scalability of the ansatz and shown that we can affect various reductions in a practical fashion to demonstrate $\mathcal{O}(N^{4})$ scaling on medium to large ab initio systems.
We are working on further improvements in the accuracy and efficiency of the ansatz, as well as taking advantage of the benefits of working in a second quantized formalism to integrate with multiscale methods and quantum embedding methodologies to provide a practical route in the modelling of truly extended systems within this CPD backflow framework [86, 86]. These techniques could also be integrated within the ‘hidden fermion’ model of correlated states as an alternative parameterization of the correlations [36, 37]. It is also natural to ask whether alternative tensor factorization techniques could be applied within the context of describing second-quantized backflow correlations. These could naturally be fitted into the framework described above, and will also be explored in the future.
Data availability
All the results presented in this work can be fully reproduced with the publicly available source code and input configurations. Energies and model parameters are further available upon request to the corresponding author.
@@
The code for this project is implemented as part of the publicly available GPSKet plugin for NetKet [63, 64]. It is made available, together with configurations files to reproduce the figures in the paper, at https://github.com/BoothGroup/GPSKet/tree/master/scripts/cpd-backflow.